OpenStudy (idealist10):

Find the equations of the trajectories of y"+y^3=0.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Do you know how to draw up the phase portrait?

OpenStudy (idealist10):

No.

OpenStudy (idealist10):

The answer is v^2+y^4/4=C I guess that p(y)=y^3, integrate p(y) to obtain P(y), P(y)=y^4/4 But I don't know how to get v^2.

OpenStudy (idealist10):

y"+p(y)=0 v(dv/dy)+p(y)=0 integrate v^2/2+P(y)=C

OpenStudy (anonymous):

I can't say I'm at all familiar with this method :/

OpenStudy (idealist10):

Yeah, the answer is pretty weird. If following the formula above, the answer should be v^2/2+y^4/4=C, but the answer in the book says v^2+y^4/4=C.

OpenStudy (anonymous):

So are we assuming \(v\) is the derivative of \(y\)?

OpenStudy (idealist10):

Well, I guess so. According to the textbook, it says y"=v(dv/dy).

OpenStudy (anonymous):

Or is \(v\) a function of \(y\)?

OpenStudy (anonymous):

Assuming \(v=v(y)\), differentiating the given solution with respect to \(y\), you have \[\frac{d}{dy}\left[v^2+\frac{y^4}{4}\right]=0~~\implies~~2v\frac{dv}{dy}+y^3=0\] which would suggest that \(y''=2v\dfrac{dv}{dy}\).

OpenStudy (idealist10):

So from y"+y^3=0, how would you do this problem?

OpenStudy (anonymous):

One thing I might try would be to multiply both sides by the first derivative: \[y'y''+y^3y'=0\] Then notice that reverse chain rule gives \[\begin{align*}\frac{d}{dt}\left[\frac{(y')^2}{2}\right]+\frac{d}{dt}\left[\frac{y^4}{4}\right]&=0\\\\ \frac{1}{2}(y')^2+\frac{1}{4}y^4&=C\end{align*}\] But past this point, I'm not sure how this could be solved...

OpenStudy (perl):

Quote `Yeah, the answer is pretty weird. If following the formula above, the answer should be v^2/2+y^4/4=C, but the answer in the book says v^2+y^4/4=C` Your book has a small typo, it should be v^2/2 + y^4/4 = C This will give you: (dy/dx)^2/2 + y^4/4 = C See siths and giggles last tell how to get that answer using reverse chain rule . Books sometimes make typos :)

OpenStudy (idealist10):

Thank you @SithsAndGiggles and @perl !

OpenStudy (perl):

solving it further is actually pretty difficult. it can be a fun challenge problem. First solve for dy/dx (dy/dx)^2/2 + y^4/4 = C (dy/dx)^2/2 = C - y^4/4 (dy/dx)^2/2 = 2C - y^4/2 dy/dx = + / sqrt (2C - y^4/2 dy / sqrt( 2C - y^4/2) = +/- dx You can integrate both sides, and you get what is known as Jacobi Elliptical integral